3 gm of oxide of metal is converted to chloride and it gives 5 gms of chloride.. Equivalent weight of metal

Grade:12th pass

1 Answers

Vikas TU
6 years ago
As we realize that Oxidation condition of oxygen is - 2 
Expect metal O.S as +1 or +2 in two cases with its nuclear wt. as X 
case 1: 
With respect to +1 express, its: M20 + 2Hcl - > 2Mcl + H2O 
2x+16 x+35.5 
in this way, it ought to be 3/(2x+16) = 2*5/(x+35.5) where x is negative, so its impractical to metal be in +1 state 
case 2: 
for +1 express, its: M0 + 2Hcl - > Mcl2 + H2O 
x+16 x+72 
along these lines, it shld be 3/(x+16) = 5/(x+72) 
which gives x = 66.5 
along these lines, the metal is Zinc 
its eq. wt = 33.25